how many quarts of pure antifreeze must be added to 2 quarts of 20% antifreeze solution to obtain a 40% antifreeze solution?
Answer by lwsshak3(11628)

how many quarts of pure antifreeze must be added to 2 quarts of 20% antifreeze solution to obtain a 40% antifreeze solution?
***
let x=amount of pure antifreeze to be added
20%*2+100%x=40%(2+x)
.4+x=.8+.4x
0.6x=0.4
x=.4/.6=2/3
amount of pure antifreeze to be added=2/3 quart